OpenStudy (anonymous):

Please help! May someone please explain step by step how to solve? My parents said I have about ten minutes to tell them the answer or they'll take my phone >.< I also want to understand though so please explain why it is the answer ♥ Find the quotient \[\frac{ x^1 }{ x^2 }\] of the complex numbers. Leave answer in polar form. \[z1=\sqrt{3}\left[ \cos \frac{ 7\pi }{ 4} +1 \sin \frac{ 7\pi }{ 4 } \right]\] \[z1=\sqrt{6}\left[ \cos \frac{ 9\pi }{ 4} +1 \sin \frac{ 9\pi }{ 4 } \right]\]

OpenStudy (anonymous):

\[[z1=\sqrt{3}\left[ \cos \frac{ 7\pi }{ 4} +1 \sin \frac{ 7\pi }{ 4 } \right]\] \[z1=\sqrt{6}\left[ \cos \frac{ 9\pi }{ 4} +1 \sin \frac{ 9\pi }{ 4 } \right]\]

OpenStudy (anonymous):

sorry...the 1 is actually an imaginary number (i) >.< I don't have my glasses...

geerky42 (geerky42):

:( so they took your phone?

geerky42 (geerky42):

HINT: \(\Large r\left[\cos(x)+i\sin (x)\right] = re^{ix}\)

geerky42 (geerky42):

So for \(\Large\dfrac{z_1}{z_2}\), you have \(\Large \dfrac{r_1 e^{ix_1}}{r_2e^{ix_2}} = \dfrac{r_1}{r_2}e^{i(x_1-x_2)} = \dfrac{r_1}{r_2}[\cos(x_1-x_2)+i\sin(x_1-x_2)]\) sorry if it looks complicated, hope it helps and you will get your phone back D:

geerky42 (geerky42):

@AngelCriner

OpenStudy (anonymous):

yes they did /.\ and I'm just seeing this I'm so sorry, thank you for explaining still! ♥